OpenStudy (liljj1421):

what is y=3(x+1)^2+9 in standard form?

OpenStudy (liljj1421):

sorry i was trying to do my other problems

OpenStudy (giannabaroni):

you're good(:

OpenStudy (liljj1421):

wouldn't you get x^2 and 1

OpenStudy (giannabaroni):

Yes(: so now what would you do to simplify this equation?

OpenStudy (liljj1421):

opps sorry idk why i put +9 it is -7

OpenStudy (liljj1421):

so would you add the +1 to the -7?

OpenStudy (photon336):

\[3(x+1)^{2} + 9= 3(x+1)(x+1) + 9\]

OpenStudy (giannabaroni):

shoot yes @Photon336 is right

OpenStudy (giannabaroni):

disregard what I said

OpenStudy (photon336):

yeah, it's easy to get tripped up on that

OpenStudy (calculusxy):

so you would actually get \[3(x+1)(x+1) = 3(x^2 + 1x + 2x + 1) = \color{green}{3(x^2 + 2x + 1)}\]

OpenStudy (calculusxy):

sorry i meant to say 3(x^2 + 1x + 1x + 1)

OpenStudy (liljj1421):

im so confused rn

OpenStudy (liljj1421):

sorry

OpenStudy (photon336):

first you need to foil (x+1)(x+1)

OpenStudy (giannabaroni):

Haha okay so (x+1)^2 is also (x+1)(x+1)

OpenStudy (giannabaroni):

so it's \[y=3(x+1)(x+1)-7\]

OpenStudy (giannabaroni):

and then foil like @Photon336 said earlier

OpenStudy (liljj1421):

what is foil?

OpenStudy (giannabaroni):

Okay so first you multiply the (F) first terms of each expression. in this case, it would be x and x. Then you multiply the (O) Outer two terms of the binomial which in this case is x and 1. After that, you do the same for the (I) Inner two terms, x and 1, and the (L) last two terms. Once you find what you got for F, O, I, and L, you add those together and simplify.

OpenStudy (giannabaroni):

@liljj1421 come back!(:

OpenStudy (giannabaroni):